# How Good is Your Wikipedia?

### v1

@@ 3.2 Heuristic Filtering
where: $S_{ij}$ is the total score for class $i$ in article $j$, $M_{ijk}$ is the value of metric $k$ in class $i$ for article $j$, $M_{min}$ is the minimum possible score for metric $k$, $M_{max}$ is the maximum possible score for metric $k$, and $K$ is the total number of metrics.
@@
An important aspect of heuristic filters is that they rely on fixed thresholds in order to discriminate between high and low quality data. In much previous work, thresholds are typically user-defined and based on background assumptions of how ‘quality’ data should be distributed. However, doing so is not possible in our case, as such assumptions are often language-specific and cannot reliably generalise in a cross-lingual fashion. Thus, we implement an automatic thresholding method that is language-agnostic and distribution-specific. We determine thresholds for filtering Wikipedia data by comparing kernel density estimates of the lowest values and a random sample of values within each metric’s distribution. The procedure is detailed in Appendix C. Doing so allows us to dynamically set thresholds across metrics and Wikipedias. This, we find, filters fewer documents than the English-based Gopher thresholds across many languages.
4 Wikipedia Quality Tiers
Figure 2: Division of Wikipedia based on a clustering on fraction of documents removed and fraction of characters removed from global filtering. The blue cluster corresponds to Tier 1, orange to Tier 2, green to Tier 3 and red to Tier 4. Each point is weighed by the article count of the unfiltered Wikipedia.
@@ 4 Wikipedia Quality Tiers
Applying the primary filtering step to Wikipedia affects some languages more than others. Thus, we categorise non-English Wikipedias into 4 tiers: Tier 1, Tier 2, Tier 3 and Tier 4 based on how much of the Wikipedia is removed during primary filtering. We perform $k$-means clustering for all the languages where each language is represented by the percentage of documents remaining and the percentage of characters remaining in the Wikipedia after primary filtering. This is done to decorrelate quality and size. Based on several iterations of the clustering algorithm and manual inspections, we arrive at the categorisation illustrated in Figure 2. A list of all Wikipedias and their respective quality tiers is in Appendix D.
To put our categorisations into perspective, we compare them to the taxonomy of languages introduced by Joshi et al. (2020). They divide the world’s languages into 6 categories based on the amount of resources and labelled and unlabelled data (including Wikipedia) available: the Left Behinds, the Scraping Bys, the Hopefuls, the Rising Stars, the Underdogs and the Winners. A comparison of these categories and our quality tiers is in Figure 3.
@@
The Joshi et al. (2020) taxonomy has a clear correlation with the size of the Wikipedia and its categorisation, making this discussion a comparison of size and quality. We find that all Winners, unsurprisingly, end up in Tier 1 and Tier 2. However, most Scraping Bys and Left Behinds do as well, indicating that many languages, despite the shortage of NLP resources and smaller Wikipedias, have dedicated communities that are involved in the creation of good quality Wikipedias. There are also many Scraping Bys and Left Behinds that are in Tier 3 and Tier 4, such as Yorùbá and Twi, with serious quality issues that can be harmful to low-resource NLP. Lastly, we also observe some Underdogs and Rising Stars in Tier 3 (Swedish, Vietnamese, Serbian and Basque) and Tier 4 (Cebuano) categories, demonstrating that more availability of resources does not guarantee good quality. Thus, the NLP developer must be aware of such quality issues when using Wikipedia.
Heuristics as a Measure of Quality
The distribution of each Wikipedia for heuristic metrics defined in §3.2 and listed in Table 1 can be an indication of the general quality of the articles in that Wikipedia. The distributions of 8 Wikipedias, 2 from each tier on metrics from classes absolute, relative and entropy, defined in Table 1, are in Figure 4.
@@
The absolute distribution shows that there is a discrepancy in the distribution of characters to articles in Wikipedia, especially for the bottom three quality tiers. Most Wikipedias have a large head of very low scoring articles (between 0 and 0.1 out of 3) and a long tail of higher scoring articles. Yorùbá, for example, has more than 50% of its character count in just 10% of articles. The relative distribution also echoes this. Most Wikipedias have a spike at 2.0, the maximum possible score for that class, assigned to articles that do not have repeating words or trigrams. This typically correlates to very short, often one-line articles. Thus, the relative class favours shorter articles, with very short articles scoring the highest, and the rest of the Wikipedia following a more normal distribution. The absolute class, on the other hand, favours longer articles. entropy metrics favour a more normal distribution for higher quality Wikipedias, in Tier 1, and the bottom tiers with lower quality articles are more skewed.
@@ 5 Downstream Evaluation
In this section we evaluate the practical implications of applying the data filters from §3 for languages in the different tiers from §4. We evaluate the effect of data pruning in two scenarios: monolingual pretraining and language adaption of multilingual models.

### v2

@@ 3.2 Heuristic Filtering
where: $S_{ij}$ is the total score for class $i$ in article $j$, $M_{ijk}$ is the value of metric $k$ in class $i$ for article $j$, $K$ is the total number of metrics, and $M_{min}$ and $M_{max}$ are the minimum and maximum possible score for metric $k$, respectively.
@@
An important aspect of heuristic filters is that they rely on fixed thresholds in order to discriminate between high and low quality data. In much previous work, thresholds are typically user-defined and based on background assumptions of how ‘quality’ data should be distributed (Rae et al., 2021; Marion et al., 2023; Raffel et al., 2023; Penedo et al., 2024). However, doing so is not possible in our case, as such assumptions are often language-specific and cannot reliably generalise in a cross-lingual fashion. Thus, we implement an automatic thresholding method that is language-agnostic and distribution-specific. This procedure is inspired by the heuristic filter tuning process of (Penedo et al., 2024), who manually inspect metric histograms for high and low quality proxy datasets (locally and globally deduplicated common crawl snapshots), and select a threshold according to where values in the low quality histogram exceed those in its high quality counterpart. Though effective, this approach does not apply to our use case in that 1) we do not have access to low or high quality proxy datasets, and 2) we operate in a multilingual space, where separate thresholds must be selected per language. Instead, we work with the heuristic metric distribution directly, taking $n$ documents from the tail (or head) to serve as a sample of potential outliers. We compare this with a random sample of $n$ documents from the overall distribution by estimating the kernel density (KDE) of both samples, which is ultimately used for determining the threshold. More details on this procedure are detailed in Appendix B. The outlined method allows us to dynamically set thresholds across metrics and Wikipedias, which, we find, filters fewer documents than the English-based Gopher thresholds across many languages.
4 Characterizing Quality
In this section, we attempt to shed light on how our chosen filtering strategies relate to a generalized notion of dataset quality, as it pertains to Wikipedia.
4.1 Establishing Quality Tiers
@@
In applying the primary filtering procedure to Wikipedia, we find that it affects certain languages much more than others. As such, we attempt to categorise all non-English Wikipedias into 4 tiers: Tier 1, Tier 2, Tier 3 and Tier 4 based on how much of the Wikipedia is removed during primary filtering. To do so, we perform $k$-means clustering, where each Wikipedia is represented by the percentage of documents and characters remaining after primary filtering. Based on several iterations of the clustering algorithm and manual inspections, we arrive at the taxonomy illustrated in Figure 2. A list of all Wikipedias and their respective quality tiers can be found in Table 16.
Figure 2: Division of Wikipedia based on a clustering on fraction of documents removed and fraction of characters removed from primary filtering. The blue cluster corresponds to Tier 1, orange to Tier 2, green to Tier 3 and red to Tier 4. Each point is weighed by the article count of the unfiltered Wikipedia.
@@
In order to put our quality taxonomy into perspective, we compare it to the taxonomy of languages introduced by Joshi et al. (2020), who divide the world’s languages into 6 categories based on the amount of labelled and unlabelled data (including Wikipedia) available for them. The categories they employ are as follows: the Left Behinds, the Scraping Bys, the Hopefuls, the Rising Stars, the Underdogs and the Winners. Figure 3 displays a comparison of these categories and our quality tiers.
Joshi et al. (2020)’s taxonomy correlates strongly with Wikipedia size, yet it yields limited alignment with our quality taxonomy. We interpret this as indicating that quality does not follow directly from resources. Indeed, while we observe that most Winners, unsurprisingly, end up in Tier 1, Arabic lands in Tier 2. Upon a cursory inspection, we find that 15% of the original Arabic Wikipedia is removed via MinHash deduplication, indicating that it contains a sizable volume of potentially templatic or bot-generated text. Going further, we find that 113 of the lowest resourced languages classified as Scraping Bys and Left Behinds (35% of Wikipedia in total) are likewise categorized as having Tier 1 Wikipedias, implying the involvement of dedicated moderation communities. Conversely, we also observe many Scraping Bys and Left Behinds languages falling into the Tier 3 and Tier 4 categories, such as Yorùbá and Twi, indicating that they contain serious quality issues that can be harmful to low-resource NLP. Lastly, we note that various higher-resourced Underdogs and Rising Stars languages are placed in Tier 3 (Swedish, Vietnamese, Serbian and Basque) and Tier 4 (Cebuano). Such Wikipedias tend to be particularly affected by deduplication: for example, Swedish loses approximately 16% and 18% of its documents to exact-match and MinHash, respectively. We interpret such findings to indicate that sheer data volume can be deceiving, and stress that NLP practitioners should be aware of such issues when utilizing Wikipedia as a resource.
4.2 Comparison to Depth+
In its official statistics, Wikimedia provides a Depth metric (alternatively, editing depth), designed to serve as a proxy measure for a given Wikipedia’s ‘collaborative quality’. [^8] https://meta.wikimedia.org/wiki/Wikipedia_article_depth However, Alshahrani et al. (2023) caution against utilizing it for such purposes. Specifically, they discuss the metric’s sensitivity to bot content (articles and edits) and excessive editing (‘edit wars’), as well as its inability to consolidate broader user activity and account for lower-resourced Wikipedias. As such, they propose a modified metric, Depth+:
$\mathrm{Depth}^{+}=\mathrm{Editors}\cdot\frac{\mathrm{Edits}}{\mathrm{Total}}\cdot\frac{\mathrm{Articles}}{\mathrm{Non-Articles}}$ (2)
Alshahrani et al. (2023) show that Depth+ is able to adequately account for bot-generated content, correctly de-ranking Wikipedias such as Vietnamese, Serbo-Croatian, and Cebuano. Conversely, they demonstrate that Depth+ likewise includes several low-resource Wikipedias, such as Cree, Tigre, and Bangla towards the top of the ranking, highlighting their users’ supposedly high degree of collaboration.
We find a significant positive correlation between our primary filtering procedure (measured as the percentage of documents retained) with Depth+ ($\rho=0.25,p<0.001$). Thus, there are similar notions of quality captured by both metrics. However, approximately half of all Wikipedias yield Depth+ values of less than 0.3, a majority very close to 0. The rest have a significantly higher Depth+ score (excluding English, German is highest with 36.04) resulting in many outliers that skew the distribution. We surmise that this is because the number of edits is disparagingly different across Wikipedia, effectively privileging Wikipedias where active users comprise a larger percentage of total users — a proxy for ‘collaborative quality’. However, this does not shed much insight on the quality of individual articles, written by one-time, potentially domain-expert authors.
4.3 Bot generation
We have observed that our primary filters remove a sizeable volume of articles from Wikipedias known to contain proliferate bot-generated content. To better understand if we are indeed filtering out such content, we refer to Alshahrani et al. (2023)’s estimations of bot-generated articles across Wikipedia. Indeed, we find that there is a strong negative correlation ($\rho=-0.56,p<0.001$) between the percentage of bot articles in a given Wikipedia and the percentage of articles remaining after primary filtering. We consider this to be an encouraging finding, provided that our primary filters are simple, pre-defined processes that state obvious assumptions about our ideal data distribution. Additionally, we do not rely on specialized bot-detection models trained on other distributions.
Figure 4: Percentage of total docs removed by MinHash and ratio of bot articles to total articles per Wikipedia.
Going further, we hypothesize that MinHash deduplication plays a large role in removing bot articles, given its ability fuzzy-match strings across documents. Here, we observe that the percentage of documents filtered out by MinHash yields an even higher correlation with bot article ratio ($\rho=0.63,p<0.001$) (Figure 4), indicating that this process is an adequate proxy for bot article removal in Wikipedia. Unsurprisingly, the Wikipedias with the highest percentage of bot content according to Alshahrani et al. (2023) — Cebuano (99%), Waray (90%), Swedish (68%) — tend to have much (if not most) of their content removed by MinHash.
4.4 Heuristics as Quality Measures
So far, we have largely focused on the effects of primary filters and their association with other recognized quality metrics. Here, explore how heuristic filters might likewise aid in our understanding of ‘quality’ as it is distributed across Wikipedia. To do so, we plot the distributions of each of the three heuristic classes defined in Table 1 across 8 Wikipedias - 2 from each tier in our quality taxonomy (Figure 5).
Immediately, we can observe that each heuristic class seems to possess its own distinct distributional type. For example, the absolute metrics tend to resemble a log-normal distribution, typically featuring a very short head and an exceedingly long tail. Intuitively, this makes sense, given that very short articles (e.g. less than ten characters) are likely to be rare, and that longer documents (but not too long) will be encountered with increasing frequency. Here, we can observe that both Hausa and Igbo (Tier 1) follow the expected pattern, featuring a short head and a gradually diminishing tail of long documents. Conversely, Waray and Minangkabau exhibit sharp, peaking heads with virtually no tail to follow them, indicating that these Wikipedias contain predominantly very short articles.
In contrast to the absolute metrics, the relative value distributions tend to be left-skewed, with the head comprising values higher than 0.5999Note that we plot combined and normalized metrics in the figure; each individual metric is bounded on $(0,1)$.. This is likewise expected, as low-quality articles for these metrics are likely to yield values closer to 0, indicating repetitive language use — for example, an article returning $\frac{3}{30}=0.1$ would feature one of three distinct tokens thirty times. Interestingly, this phenomenon is generally observed for all 8 languages, with the of exception of Amharic, which features a narrow, short mode around 0.5 — possibly indicating a cluster of low quality data.
Lastly, the entropy distributions tend to be somewhat normal in shape, which is indeed the desired outcome for this metric, signifying that documents follow predictable n-gram patterns. Generally, these metrics are somewhat related to the relative class, in that lower values will tend to correspond to highly predictable data. However, while higher-valued relative metrics are typically desirable, coinciding with varied lexical usage, higher-valued entropy values correspond instead to completely unpredictable — and potentially undecipherable — text. In the latter case, this could manifest by HTML or javascript leakage or poor page formatting. Interestingly, three of the four lower-tiered languages (Yoruba, Waray, Minangkabau) appear to exhibit a higher concentration of lower-entropy documents, indicating that they include more predictable, potentially bot-generated text. Conversely, Hausa and Igbo are left-skewed, implying that their articles feature more creative language use.
@@ 5 Downstream Evaluation
Having demonstrated that our filtering procedures align with various qualitative notions of ‘quality’, we now turn to a quantitative evaluation focused on downstream NLP tasks. Specifically, we evaluate the effect of data pruning in three scenarios: monolingual pretraining, language adaptation, and multilingual pretraining (see Figure 6).